OpenStudy (anonymous):

If a regular polygon has exterior angles that measure 60 degrees each, how many sides does the polygon have?

jimthompson5910 (jim_thompson5910):

n = 360/E n = 360/60 n = 6 so there are 6 sides and this polygon is a hexagon

OpenStudy (anonymous):

Thank You

jimthompson5910 (jim_thompson5910):

you're welcome